OpenStudy (anonymous):

whats the y intercept and slope for y=15x+15

geerky42 (geerky42):

You are given slope-intercept equation. \(y = mx+b\) m is slope and b is y-intercept. Do you know what m and b is?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hello?

geerky42 (geerky42):

OK so what are they?

OpenStudy (anonymous):

its y=mx+b

OpenStudy (anonymous):

heres the full problemThe entrance fee to a national park is $15. A campsite fee is $15 per night. The total cost y for a camping for x nights can be represented by the equation y=15x+15. graph the eqaution and find slope

geerky42 (geerky42):

Can you graph it? Slope is 15.

OpenStudy (anonymous):

yea its graphed it would be 0=15 and 3=60

geerky42 (geerky42):

Yeah, just find two points, then draw line through points. So is everything clear?

OpenStudy (anonymous):

so the y intercept is the same thing as the eqaution

OpenStudy (anonymous):

i have another question similiar to this one.

OpenStudy (anonymous):

write an equation of a line in slope intercept form with the given slope and y intercept. slope:4 y; y-inetrcept:1

geerky42 (geerky42):

general line equation is y = mx+b; m is slope and b is y-intercept So since you are given that m=4 and b=1, you can replace m and b to 4 and 1 respectively. So you have \(y = 4x+1\) Is that clear?

OpenStudy (anonymous):

oh ok so. slope: 1/2 ; y- intercept: = - 3. would be y=1/2x+-3

geerky42 (geerky42):

Right, or \(y = \frac{1}{2}x-3\)

geerky42 (geerky42):

adding -3 is same as subtracting 3.

OpenStudy (anonymous):

ok thanks and the graph one would have slope of 15 and y intercept of y=15x+15 same as the question?

geerky42 (geerky42):

Well, slope is 15 and y-intercept is 15, yeah.

OpenStudy (anonymous):